# Improving LLM First-Token Predictions in Multiple-Choice Question Answering via Prefilling Attack

### v1

@@ A.2 Open-Ended Generation
Prompting and Classification Setup. We leverage GPT-3.5-Turbo and Llama-3.1-70B-Instruct to classify the open-ended generations produced by the models in response to benchmark questions. To generate these responses, we follow the same prompting procedure used for the FTP method: the model is presented with the question and all possible answer options, and in this case, it is allowed to generate freely until the EOS token.
@@
To convert each open-ended response into a symbolic answer label, we prompt GPT-3.5-Turbo and Llama-3.1-70B-Instruct with the following instruction:
@@
and response is the text generated by the evaluated model. In both cases, the predicted classification is consistently returned either as the bare letter (e.g., ‘B’) or the letter followed by parenthesis (e.g., ‘B)’). We then compare the predicted letter with the ground-truth answer to compute the final accuracy scores.
Expanded Comparison with Llama-Based Classification. In addition to the analysis presented in Section 4.3, in Figure 5 we use Llama-3.1-70B-Instruct to classify the open-ended generations produced by all evaluated models across all benchmark datasets. We opt for Llama-3.1-70B for this evaluation because it is openly accessible and facilitates reproducibility.
As shown, the accuracies obtained from Llama-classified generations are on par with those of the FTP approach enhanced with prefilling across all datasets and models, indicating that our prefilling strategy is a robust and effective evaluation strategy. The only case where we observe noticeably lower results compared to open-ended classification is MathQA. The reduced accuracy of both FTP and FTP with prefilling on this dataset suggests that more advanced reasoning is required, and that prefilling alone cannot compensate for this limitation.
@@
Table 7: Full-vocabulary first-token evaluation on MMLU, OBQA, SIQA, and SCIQ, for all of our tested models. The model must predict a valid and correct token as its very first output. Prefilling improves both validity and full-vocabulary accuracy.
A.3 Full Results on Full-Vocabulary Evaluation
Table 7 complements the partial results in Table 2 by reporting full-vocabulary accuracy and First Token Validity Rate (FTVR) for all models and benchmarks. While most models benefit from prefilling, this full breakdown reveals particularly brittle behavior in the Gemma models: both Gemma-7B and Gemma-2-9B consistently produce 0% full-vocabulary accuracy and 0% FTVR under standard FTP. This failure is due to their tendency to always begin responses with the token ‘The’, forming preambles like ‘The correct answer is’ instead of immediately emitting a symbolic option. As a result, their top-1 token never matches the valid label set when decoding is unconstrained. The prefilling strategy successfully overrides this default behavior, enabling proper symbolic alignment and resulting in large performance gains across different MCQA benchmarks.

### v2

@@ Appendix 0.B Open-Ended Generation
Prompting and Classification Setup. We leverage GPT-3.5-Turbo, Llama-3.1-70B-Instruct, and xFinder-Qwen to classify the open-ended generations produced by the models in response to benchmark questions. To generate these responses, we follow the same prompting procedure used for the FTP method: the model is presented with the question and all possible answer options, and in this case, it is allowed to generate freely until the EOS token.
@@
To convert each open-ended response into a symbolic answer label, we prompt GPT-3.5, Llama-3.1, and xFinder with the following instruction:
@@
and response is the text generated by the evaluated model. In all cases, the predicted classification is consistently returned either as the bare letter (e.g., ‘B’) or the letter followed by parenthesis (e.g., ‘B)’). We then compare the predicted letter with the ground-truth answer to compute the final accuracy.
Expanded Comparison with Llama- and xFinder-Based Classification. In addition to the analysis presented in Sec. 4.3 of the main paper, we evaluate the open-ended generations produced by all models across the benchmark datasets. The results using Llama-3.1-70B-Instruct are reported in Fig. 5, while the results using the xFinder-Qwen classifier are reported in Fig. 6. Llama-3.1-70B-Instruct is chosen for its accessibility and reproducibility, whereas xFinder-Qwen is chosen because it is optimized for accurately identifying and extracting answers from LLM-generated text.
As shown, the accuracies obtained from Llama-classified generations are on par with those of the FTP approach enhanced with prefilling across all datasets and models, indicating that our prefilling strategy is a robust and effective evaluation strategy. The only cases where we observe noticeably lower results compared to open-ended classification are MathQA and ARC-C. The reduced accuracy of both FTP and FTP with prefilling on these datasets suggests that more advanced reasoning is required, and that prefilling alone cannot compensate for this limitation.
A similar trend is observed with xFinder-classified results: in most cases, FTP with prefilling achieves higher accuracies, while for OpenBookQA and Moral Stories, FTP with prefilling and open-ended generation evaluated via xFinder perform roughly equally. For MathQA, results are more mixed, with open-ended generation sometimes providing an advantage, but not consistently across all models. Nonetheless, MathQA results highlight our main message: even when prefilling does not close the gap to open-ended reasoning in reasoning-heavy tasks, it still provides consistent gains without additional compute.
Appendix 0.C Full Results on Full-Vocabulary Evaluation
Table 6 complements the partial results in Table 2 of the main paper by reporting full-vocabulary accuracy and First Token Validity Rate (FTVR) for all models and benchmarks. While most models benefit from prefilling, this full breakdown reveals particularly brittle behavior in the Gemma models: both Gemma-7B and Gemma-2-9B consistently produce 0% full-vocabulary accuracy and 0% FTVR under standard FTP. This failure is due to their tendency to always begin responses with the token ‘The’, forming preambles like ‘The correct answer is’ instead of immediately emitting a symbolic option. As a result, their top-1 token never matches the valid label set when decoding is unconstrained. The prefilling strategy successfully overrides this default behavior, enabling proper symbolic alignment and resulting in large performance gains across different MCQA benchmarks.
Table 6: Full-vocabulary first-token evaluation on four MCQA benchmarks for all tested LLMs. The model must predict a valid and correct token as its very first output. Prefilling improves both validity and full-vocabulary accuracy.
@@ Appendix 0.C Full Results on Full-Vocabulary Evaluation
Appendix 0.D Calibration Metrics
Here, we provide detailed definitions and explanations of the calibration metrics used in Sec. 4.5 of the main paper. In particular, Adaptive Calibration Error (ACE) [31] partitions predictions into adaptive ranges such that each range contains an equal number of predictions, mitigating bias from sparsely populated bins and better reflecting calibration across all classes. The adaptive ranges are created by first sorting the predicted probabilities for each class. The sorted list is then divided into $R$ contiguous ranges such that each range contains approximately $\lfloor N/R\rfloor$ predictions, where $N$ is the total number of predictions for that class. This ensures that each range is equally populated, allowing ACE to focus on regions with sufficient data while avoiding the sparsity issues that affect fixed bins. Formally, let $R$ be the number of adaptive ranges and $K$ the number of classes. ACE is defined as
$\text{ACE}=\frac{1}{KR}\sum_{k=1}^{K}\sum_{r=1}^{R}\left|\text{acc}(r,k)-\text{conf}(r,k)\right|,$ (3)
where $\text{acc}(r,k)$ and $\text{conf}(r,k)$ denote the empirical accuracy and average predicted confidence for class $k$ in adaptive range $r$, respectively. Unlike Expected Calibration Error (ECE) [29], which only considers the probability of the predicted class for each example, ACE evaluates calibration for each class separately by including the predicted probabilities for all classes, not just the one with the highest predicted probability.
The Brier score [11] is the mean squared error between the predicted probability $p_{i}$ for the correct class and the true label $y_{i}$ across all $n$ examples:
$\text{Brier Score}=\frac{1}{n}\sum_{i=1}^{n}(p_{i}-y_{i})^{2}.$ (4)
It quantifies both the accuracy and the calibration of the probabilistic predictions, penalizing confident incorrect predictions.
Log Loss [12, 19], also known as negative log likelihood or cross-entropy loss, evaluates the probabilistic predictions by aggregating the log-probabilities assigned to the true classes. For a single example with true class $y$ and predicted probabilities $p_{k}$, it is defined as
$\text{Log Loss}(x_{i},y_{i})=-\sum_{k=1}^{K}y_{i,k}\log(p_{i,k}),$ (5)
where $K$ is the number of classes, $y_{i,k}$ is 1 if class $k$ is the true label and 0 otherwise, and $p_{i,k}$ is the predicted probability assigned to class $k$.
The overall Log Loss across a dataset of $n$ examples is then computed as the average:
$\text{Log Loss}=\frac{1}{n}\sum_{i=1}^{n}\text{Log Loss}(x_{i},y_{i}).$ (6)
This formulation penalizes predictions that assign low probability to the true class; equivalently, it severely penalizes overconfident misclassifications in which the model assigns high probability to an incorrect class. Compared to the Brier score, which applies a quadratic penalty, Log Loss is harsher on highly confident errors.
Appendix 0.E Limitations
We hope that future work will extend our analysis to broader tasks and languages, and will develop decoding strategies and prompt formats that are inherently robust to structural ambiguity and symbolic misalignment.
Figure 6: Accuracy comparison on all the benchmarks using standard FTP, FTP with prefilling, and open-ended generation with xFinder-Qwen as classifier. Again, FTP with prefilling consistently outperforms standard FTP and often outperforms or is on par with the more computationally expensive open-ended generation approach.